The length of a rectangle is 10 more than twice the width. The area of the rectangle is 672 square inches. find the length and the width. (Remember the dimensions cannot be negative.)
Answer by Boreal(15235)

W=x
L=2x+10
A=(2x+10)*x=2x^2+10x=672
2x^2+10x-672=0
x^2+5x-336=0
(x+21)(x-16)=0
x=16 inches width
2x+10=42 inches length.
That area is 672 in^2.